Martina leaves home at 9 A.M., bicycling at a rate of 24 mi/h. Two hours later, John leaves, driving at the rate of 48 mi/h. At what time will John catch up with Martina?
Answer by ankor@dixie-net.com(22740)

Martina leaves home at 9 A.M., bicycling at a rate of 24 mi/h. Two hours later, John leaves, driving at the rate of 48 mi/h. At what time will John catch up with Martina?
:
When J catches up with M, they will have traveled the same distance, make a distance equation from that fact:
:
Let t = time required for J to catch up with M
Then
(t+2) = time that M will be on the road when this happens
:
Dist = speed * time
:
J's Dist = M's dist
48t = 24(t+2)
48t = 24t + 48
48t - 24t = 48
t = 48/24
t = 2 hrs
:
J left at 11 so at 1 PM J will catch up with M
:
:
Check solution by finding if the distances are equal
48(2) = 24(2+2)